Circular No. 6610

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970402 M. Feroci, E. Costa, and L. Piro, Istituto Astrofisica Spaziale, CNR,

Frascati; F. Frontera and G. Zavattini, Universitá di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; G. Celidonio, G. Gandolfi, and A. Coletta, BeppoSAX, Rome; and J. Heise, and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, report: “The BeppoSAX Gamma Ray Burst Monitor (GRBM) was triggered on Apr. 2.930 UT. A quick-look analysis of the GRBM light curve shows a rather weak, highly structured γ-ray burst (GRB) lasting > 100 s. The intensity at its maximum is ∼ 300 counts/s. The γ-ray burst was detected by the Wide Field Camera (WFC) number 1 with a similar time profile and a maximum countrate of ∼ 70 counts/s. Taking into account the off-axis position, the peak flux corresponds to ∼ 0.46 Crab. From the WFC image, we have derived a preliminary position of α = 14^{h}50^{m}24^{s}, δ = −69^{o}18′.0 (equinox 2000.0), with a 12′ error radius. A refined position will result from the off-line analysis. A BeppoSAX TOO pointing of the Narrow Field Instruments was performed starting on Apr. 3.270, just 8 hr after the GRB trigger time. Observations in all bands are solicited.”

J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands; E. Costa, Istituto di Astrofisica Spaziale, CNR, Frascati; and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report on behalf of the BeppoSAX-WFC team: “GRB 970402 was localized in the field of Wide Field Camera 1 using the known positions of five other x-ray sources in the same field (e.g., Cir X-1, 4U 1254-63, and Cen X-3) as verification. The position of GRB 970402 is α = 14^{h}50^{m}16^{s}, δ = −69^{o}19′.9 (equinox 2000.0), with an error-circle radius of 3′.”

SUPERNOVA 1997bm IN IC 705 D. Balam, University of Victoria, reports that Cousins R-band im-

ages obatined with the 1.82-m Plaskett Telescope of the National Research Council of Canada on Apr. 3.29 UT (when R = 18.1 ± 0.2) yield the following position for SN 1997bm: α = 11^{h}32^{m}57^{s}.13, δ = +50^{o}14′43′′.6 (equinox 2000.0), which is 7′′.8 east and 12′′.9 north of the galaxy’s center. A nearby star (R = 17.2 ± 0.2) has position end figures 51^{s}.79, 13′11′′.2. Several long V -band images (300 and 600 s) suggest that the supernova is rather red.

1997 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green